OpenStudy (anonymous):

If k=x+3/x-3, and k^2-k=12, find x. (show work)

OpenStudy (mathmale):

Hint: the 2nd equation can be written in the form of a quadratic equation. Do so, please, and solve that equation for the 2 different values of k. Then subtitute those k values, one at a time, into the first equation.

OpenStudy (anonymous):

so then the quadratic equation is: K^2-K-12=0 correct?

OpenStudy (mathmale):

Yes. Now please solve that eq'n for k.

OpenStudy (anonymous):

can you give me another hint, this is helping me a lot. what is the value of K

OpenStudy (mathmale):

K^2-K-12=0 can be factored: (k-4)( ? ). Can you finish that?

OpenStudy (anonymous):

perfect, so then (K-4) and (K+3) because multipled -4 * 3 = -12 and added = -1 this is very helpful.

OpenStudy (anonymous):

so then k1 = -3 and k2 = 4

OpenStudy (anonymous):

but then do i substitute back into the equation?